# Mechanistic Interpretability of Large-Scale Counting in LLMs through a System-2 Strategy

### v1

@@ 7 Conclusion
This paper showed that the failure of large language models on large-scale counting tasks arises from reliance on System-1–like implicit counting mechanisms with limited capacity (kahneman2011thinking; zhang2024countingabilitytokenization). This problem reflects architectural constraints of transformer models rather than fundamental limits of numerical reasoning. We introduced a simple System-2 test-time strategy that decomposes large counting problems into smaller sub-tasks and aggregates their results, enabling accurate counting over long contexts without modifying model parameters or training procedures.
In addition to behavioral improvements, we provided a mechanistic explanation of how this strategy operates internally. Using attention analysis and causal mediation methods (hasani2025countscope; heimersheim2024activationpatching), we showed that partition-level counts are encoded at boundary tokens, transferred through specific attention pathways to intermediate reasoning steps, and aggregated in middle-to-late layers to form the final answer. Intermediate reasoning tokens play a crucial role in this process, mediating the flow of numerical information from input partitions to the final output.
In principle, this procedure removes fixed upper bound on countable size, as long as each sub-problem remains within the model’s reliable regime. While our experiments focus on counting, the same analysis framework applies to other reasoning tasks where implicit representations saturate and explicit decomposition enables correct behavior. This study highlights how structured test-time strategies can reveal and extend the computational abilities of existing models, offering a path toward both improved performance and deeper interpretability. Future research could explore further applications of this strategy to complex reasoning tasks in language and beyond.
@@
This work focuses on a narrow counting task with limited object diversity, using simple repeated nouns in synthetic contexts. The approach depends on structured prompts and assumes prior knowledge of the model’s reliable counting range, which may limit generalization. In addition, the effects of tokenization and number representation are not isolated and may independently influence the results.
@@ Appendix A Appendix
A.1 Details of the Behavioral Setup
@@
A.2 Attention Analysis
This section presents detailed methodological information regarding the attention analyses conducted in this study. We perform our analysis using various large language models, Qwen 2.5 7B (qwen25), Llama 3.2 8B (llama3), and Gemma 3 4B(gemma). These models differ in architectural depth and layer organization, which allows for a more comprehensive and comparative investigation of attention behaviors. The analysis is divided into two parts. In the first part, we describe the experimental setup in detail, including the selection of specific layers for analysis and the prompts used in each experiment. In the second part, we analyze the attention patterns of the generated outputs with respect to the full input sequences for both models.
A.2.1 Input Details
The exact prompt used in all experiments is shown below:
@@ A.2.1 Input Details
A.2.2 Full Attention Analysis
This section presents a full attention analysis (following the methodology in Fig. 4) for Qwen2.5 7B, Gemma 3 4B, and Llama 3.2 8B. Our goal is to explore a little more of the layers in which attention most strongly concentrates on the tokens relevant to partition prediction, and to visualize the resulting full attention patterns.
As shown in Fig. 11, we analyze attention scores across different layer ranges for each model. For Gemma 3 4B, we examine layers 21 to 23, while for Llama 3.2 8B we consider layers 13 to 18. These layer intervals are selected based on the average attention magnitude from output tokens to key input tokens (specifically, the last item and the trailing comma of the predicted partition), which consistently peak within these ranges.
Figures 12, 13, and 14 illustrate the full attention maps for Qwen2.5 7B, Llama 3.2 8B, and Gemma 3 4B, respectively. In each figure, the x-axis corresponds to output tokens and the y-axis corresponds to input tokens. Across all models, we observe a consistent pattern: the token immediately preceding the generated number of parts exhibits the highest attention weights toward the last item and the final comma of the desired partition. This behavior supports our hypothesis that models rely heavily on attention to the final item–comma structure of each part when determining and generating the number of items for each part.
@@ A.2.2 Full Attention Analysis
(b) Average attention weights across layers for Gemma 3 4B. We report attention from selected output tokens to both input and output tokens (as defined in Fig. 4(a)). Attention magnitude peaks between layers 21 and 23.
Figure 11: Layer-wise attention analysis for Llama 3.2 8B and Gemma 3 4B. For each model, we visualize the average attention from selected output tokens to salient input tokens (notably the final item and comma of the partition). These trends are used to identify the layer ranges with the strongest partition-relevant attention.
@@
Figure 13: Full attention map for Llama 3.2 8B. The x-axis corresponds to output tokens and the y-axis to input tokens. The token immediately preceding the generated number of parts exhibits the strongest attention toward the final item and trailing comma of each partition matches its own segment.
Figure 14: Full attention map for Gemma 3 4B. The x-axis corresponds to output tokens and the y-axis to input tokens. The token immediately preceding the generated partition-level count exhibits the strongest attention toward the final item and trailing comma of each partition matches its own segment.
A.3 Causal Mediation Analysis
@@
Figure 15: Average change in log probability of intermediate counts after layer masking (a) and unmasking (b) of the final item and separator of each partition. In the masking experiment, embeddings from the target layer are zero-ablated. In the unmasking experiment, embeddings of the selected tokens are zero-ablated across all layers, and only the target-layer embeddings are restored from the clean run.

### v2

@@ 7 Conclusion
This paper showed that the failure of large language models on large-scale counting tasks arises from reliance on System-1–like implicit counting mechanisms with limited capacity (Kahneman, 2011; Zhang et al., 2024). This problem reflects architectural constraints of transformer models rather than fundamental limits of numerical reasoning. We introduced a simple System-2 test-time strategy that decomposes large counting problems into smaller sub-tasks and aggregates their results, enabling accurate counting over long contexts without modifying model parameters or training procedures.
In addition to behavioral improvements, we provided a mechanistic explanation of how this strategy operates internally. Using attention analysis and causal mediation methods (Hasani et al., 2025b; Heimersheim and Nanda, 2024), we showed that partition-level counts are encoded at boundary tokens, transferred through specific attention pathways to intermediate reasoning steps, and aggregated in middle-to-late layers to form the final answer. Intermediate reasoning tokens play a crucial role in this process, mediating the flow of numerical information from input partitions to the final output.
In principle, this procedure removes a fixed upper bound on countable size, as long as each sub-problem remains within the model’s reliable regime. While our experiments focus on counting, the same analysis framework may extend to other reasoning tasks where implicit representations saturate and explicit decomposition into sub-tasks facilitates correct behavior. This study highlights how structured test-time strategies can reveal and extend the computational abilities of existing models, offering a path toward both improved performance and deeper interpretability. Future research could explore further applications of this strategy to more complex reasoning tasks in language and beyond multimodal settings.
@@
This work focuses on a text-based counting task with limited object diversity. For systematic mechanistic analysis, we use synthetic contexts with repeated nouns rather than natural free-form text. The approach depends on structured prompts and assumes prior knowledge of the model’s reliable counting range. While this requirement is mild in practice, it may limit applicability across domains. The proposed decomposition strategy is most effective for tasks that can be divided into near-independent sub-tasks (e.g., multi-step arithmetic or sequential planning). It is not expected to generalize to tasks with strong interdependencies, where subproblems cannot be cleanly separated.
@@ Appendix A Appendix
A.1 Details of the Experimental Setup
@@
Model Input Output 11–20 21–30 31–40 41–50
Gemma2-2B Unstructured w/o steps 0.01 0.00 0.00 0.00
w/ steps 0.00 0.00 0.00 0.00
Structured w/o steps 0.05 0.12 0.03 0.04
w/ steps 0.18 0.15 0.10 0.07
Gemma3-4B Unstructured w/o steps 0.91 0.45 0.04 0.08
w/ steps 0.99 0.52 0.02 0.03
Structured w/o steps 0.26 0.13 0.03 0.05
w/ steps 0.96 0.56 0.19 0.17
Table 4: Average accuracy on smaller open-source models (Gemma2-2B and Gemma3-4B) for context sizes from 11 to 50. Each model is evaluated on structured and unstructured inputs, with and without intermediate reasoning steps.
Model Input Output 11–20 21–30 31–40 41–50
Qwen2.5-Math-7B Unstructured w/o steps 0.95 0.54 0.04 0.04
w/ steps 0.94 0.46 0.02 0.05
Structured w/o steps 0.97 0.89 0.89 0.75
w/ steps 1.00 0.77 0.75 0.52
DeepSeekMath-7B Unstructured w/o steps 0.86 0.12 0.04 0.04
w/ steps 0.90 0.30 0.03 0.05
Structured w/o steps 0.40 0.21 0.03 0.01
w/ steps 0.86 0.58 0.38 0.19
Table 5: Average accuracy on math-specialized models for context sizes from 11 to 50. DeepSeekMath-7B follows the same trend as general-purpose models, where structured input with intermediate reasoning performs best. Qwen2.5-Math-7B shows stronger performance for both structured settings, suggesting improved internal aggregation of partition-level counts.
A.2 Behavioral Results on Additional Models
We extend the behavioral evaluation to two smaller open-source models, Gemma2-2B (Riviere et al., 2024) and Gemma3-4B (Kamath et al., 2025), and two math-specialized models, Qwen2.5-Math-7B (Yang et al., 2024b) and DeepSeekMath-7B (Shao et al., 2024). We use the same four settings as in the main paper: unstructured vs. structured inputs, each with short-answer generation or intermediate reasoning steps. We evaluate context sizes from 11 to 50 items, grouped into four bins.
Results for smaller models are shown in Table 4. As expected, absolute accuracy decreases with model size, especially for longer contexts. However, the main trend from the paper remains unchanged. Structured input with intermediate reasoning is the strongest setting overall and remains more robust than the other three variants as context length increases. This suggests that the proposed decomposition strategy is not limited to larger models and still provides benefits in lower-capacity regimes.
Results for math-specialized models are shown in Table 5. DeepSeekMath-7B follows the same pattern as general-purpose models, where structured input with intermediate reasoning gives the best performance. Qwen2.5-Math-7B shows a different behavior. Both structured settings perform strongly, and structured input without intermediate reasoning is often competitive with, or better than, the version with reasoning steps. This suggests that some math-tuned models may better combine partition-level counts internally without requiring explicit intermediate outputs. This behavior is also consistent with stronger out-of-context reasoning ability, where some models remain effective in the no-CoT regime (Deng et al., 2026; Yao et al., 2025).
These additional experiments support the main conclusion that external structure is highly useful for long-range counting. For most models, the strongest configuration remains structured input with intermediate reasoning. At the same time, the Qwen2.5-Math result suggests that domain-specific fine-tuning can partially change how aggregation is carried out at inference time while maintaining strong performance in the no-CoT setting.
A.3 Robustness to Tokenization and Input Structure
We evaluate the generality of our System-2 counting mechanism across different tokenization schemes and input formats. Table 6 summarizes the tokenizer types and vocabulary sizes of the models used in our main experiments. Despite these differences, the overall behavioral and mechanistic patterns are consistent, indicating that the counting strategy does not rely on a specific tokenizer or vocabulary.
Model Tokenizer Type Vocabulary Size
GPT-4o BPE $\sim$200k
Gemini 2.5 Pro SentencePiece $\sim$256k
Gemma 3 SentencePiece $\sim$262k
Llama 3 BPE $\sim$128k
Qwen 2.5 Byte-level BPE $\sim$150k
Table 6: Tokenizer types and vocabulary sizes for the evaluated models.
To further assess robustness, we evaluate two alternative input structures with different separators and instructions:
Structure A: Intermediate structures use part labels with commas separating items, e.g.,
part1: apple, apple
part2: apple, apple, apple
Structure B: Intermediate structures list items with spaces, and parts are separated by slashes, e.g.,
apple apple / apple apple apple
Table 7 shows that the same performance trends hold for Qwen2.5-7B: structured inputs combined with intermediate reasoning steps consistently outperform unstructured inputs or inputs without reasoning steps, confirming that the System-2 mechanism is robust to minor changes in input formatting. However, the superior performance of Structure A over Structure B highlights the role of input structure in overall performance.
Structure Input Output 11–20 21–30 31–40 41–50
A Unstructured w/o steps 0.30 0.11 0.05 0.00
Unstructured w/ steps 0.45 0.10 0.02 0.00
Structured w/o steps 0.38 0.24 0.16 0.10
Structured w/ steps 1.00 0.77 0.64 0.32
B Unstructured w/o steps 0.28 0.09 0.05 0.01
Unstructured w/ steps 0.28 0.10 0.04 0.01
Structured w/o steps 0.22 0.19 0.12 0.10
Structured w/ steps 0.74 0.46 0.39 0.24
Table 7: Accuracy for Qwen2.5 7B under two different input formats (A and B) with structured and unstructured inputs, with and without intermediate reasoning steps.
A.4 Attention Analysis
Input Details
The exact prompt structure is provided below:
@@ A.4 Attention Analysis
Further Attention Results
We extend the attention analysis to several large language models, including Qwen2.5-7B (Yang et al., 2024a), Llama3.2 8B (Grattafiori et al., 2024), and Gemma3-4B (Kamath et al., 2025). These models differ in architectural depth and layer organization, enabling a more comprehensive comparative analysis of attention behaviors across architectures.
As shown in Figure 12, we analyze attention scores across different layer ranges for each model. For Gemma3-4B, we examine layers 21 to 23, while for Llama3.2-8B we consider layers 13 to 18. These layer intervals are selected based on the average attention magnitude from output tokens to key input tokens (specifically, the last item and the trailing comma of the predicted partition), which consistently peak within these ranges.
Moreover, Figures 13, 14, and 15 illustrate the full attention maps for Qwen2.5-7B, Llama3.2 8B, and Gemma3-4B, respectively. In each figure, the x-axis corresponds to output tokens and the y-axis corresponds to input tokens. Across all models, we observe a consistent pattern: the token immediately preceding the generated number of parts exhibits the highest attention weights toward the last item and the final comma of the desired partition. This behavior supports our hypothesis that models rely heavily on attention to the final item–comma structure of each part when determining and generating the number of items for each part.
@@
(b) Average attention weights across layers for Gemma3-4B. We report attention from selected output tokens to both input and output tokens (as defined in Figure 4(a)). Attention magnitude peaks between layers 21 and 23.
Figure 12: Layer-wise attention analysis for Llama3.2-8B and Gemma3-4B. For each model, we visualize the average attention from selected output tokens to salient input tokens (notably the final item and comma of the partition). These trends are used to identify the layer ranges with the strongest partition-relevant attention.
@@
Figure 14: Full attention map for Llama3.2-8B. The x-axis corresponds to output tokens and the y-axis to input tokens. The token immediately preceding the generated number of parts exhibits the strongest attention toward the final item and trailing comma of each partition matches its own segment.
Figure 15: Full attention map for Gemma3-4B. The x-axis corresponds to output tokens and the y-axis to input tokens. The token immediately preceding the generated partition-level count exhibits the strongest attention toward the final item and trailing comma of each partition matches its own segment.
A.5 Causal Mediation Analysis